Grade 12General Physics

The no. of atoms of a radioactive substance of half-life T is N0 at t=0. The time necessary to decay from N0/2 atoms to N0/10 atoms wil be
  1. 5T/2
  2. T ln 5
  3. T ln(5/2)
  4. T(ln 5/ln 2)

Indrani
9 Years agoGrade 12

1 Answer

Khimraj
9 Years ago
T = (ln2)/λ and lnNo/N = λt. So time for No to No/10 will be (ln10)/λ and from No/2 to No/10 will be ln10/λ - ln2/λ = ln5/λ = T(ln5/ln2).Hope it clears. If u have doubts then please clearify. And if u like my answer then please approve answer